OpenStudy (anonymous):

This one is so hard! ): Given F(x)=5x^6 and G(x)=15x^-2 what is F(x) divided by G(x)?

OpenStudy (anonymous):

@kyusakazaki @zzr0ck3r

OpenStudy (zzr0ck3r):

\[\frac{a^b}{a^c}=a^{b-c}\]so\[\frac{5x^6}{15x^{-2}}=\frac{1}{3}x^{6-(-2)}=\frac{1}{3}x^?\]

OpenStudy (anonymous):

how do i get the answr out of that?

OpenStudy (anonymous):

is it x/3^4

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

@mathcalculus do you understand this?

OpenStudy (anonymous):

@zzr0ck3r do you know the answr?

OpenStudy (luigi0210):

We know the answer, but we are not allowed to give it to you. We are here to walk you through the process as @zzr0ck3r did

OpenStudy (anonymous):

yeah well i did not get it when he walked me through the process and im in need of the answer.

OpenStudy (zzr0ck3r):

what is x^2/x?